### v1

# Enhanced DACER Algorithm with High Diffusion Efficiency

Abstract
Due to their expressive capacity, diffusion models have shown great promise in offline RL and imitation learning. Diffusion Actor-Critic with Entropy Regulator (DACER) extended this capability to online RL by using the reverse diffusion process as a policy approximator, trained end-to-end with policy gradient methods, achieving strong performance. However, this comes at the cost of requiring many diffusion steps, which significantly hampers training efficiency, while directly reducing the steps leads to noticeable performance degradation. Critically, the lack of inference efficiency becomes a significant bottleneck for applying diffusion policies in real-time online RL settings. To improve training and inference efficiency while maintaining or even enhancing performance, we propose a Q-gradient field objective as an auxiliary optimization target to guide the denoising process at each diffusion step. Nonetheless, we observe that the independence of the Q-gradient field from the diffusion time step negatively impacts the performance of the diffusion policy. To address this, we introduce a temporal weighting mechanism that enables the model to efficiently eliminate large-scale noise in the early stages and refine actions in the later stages. Experimental results on MuJoCo benchmarks and several multimodal tasks demonstrate that the DACER2 algorithm achieves state-of-the-art performance in most MuJoCo control tasks with only five diffusion steps, while also exhibiting stronger multimodality compared to DACER.
1 Introduction
Figure 1: Diffusion steps and performance. Taking the Walker2d-v3 task as an example.
Diffusion models have recently demonstrated remarkable performance in image generation epstein2023diffusion ; zhang2023text ; miao2024training and video generation ceylan2023pix2video ; yang2024cogvideox ; wang2025lavie . Owing to their expressive capabilities, diffusion models can represent a wide range of complex distributions, rendering them highly effective for diverse generative tasks. This expressiveness also positions them as a suitable policy class for continuous control, commonly referred to as diffusion policies ren2024DPPO ; li2024learning ; lu2025makes . Moreover, their generative nature aligns naturally with offline reinforcement learning (RL) paradigms wang2022diffusion ; ada2024diffusion ; kang2023efficient .
In online RL, adopting an energy-based model as the policy equips the agent with powerful representational capabilities. Learning a policy to approximate the corresponding energy-based target distribution allows for modeling complex and multimodal action patterns without relying on restrictive parametric assumptions, especially in continuous action spaces. This enhanced expressiveness can significantly improve exploration by enabling the agent to discover and exploit diverse behavioral strategies. However, effectively approximating such an expressive soft policy presents notable challenges. A key difficulty lies in how to efficiently and accurately sample from the target distribution. While algorithms such as Soft Actor-Critic (SAC) haarnoja2018soft and Distributional Soft Actor-Critic (DSAC) duan2021distributional ; duan2023dsac aim to approximate the soft-target distribution, they generally represent the policy as a simple Gaussian. This choice is computationally efficient but falls short in modeling complex, multimodal behavior.
Recent advances have explored more expressive sampling approaches, such as diffusion models psenka2023learning ; ma2025soft ; messaoud2024s , to better approximate the Boltzmann distribution. Existing methods can be broadly categorized into two groups: score matching and policy gradient approaches. In the first group, QVPO ding2024diffusion proposes using Q-weighted imitation learning samples to improve policy learning. QSM psenka2023learning directly aligns the score functions with the gradients of the learned Q functions and uses Langevin dynamics for sampling. DIPO yang2023diffusion updates the replay buffer using action gradients and improves the performance of the policy through a diffusion loss. In the second group, DACER wang2024diffusion directly backward the gradient through the reverse diffusion process and proposes a GMM entropy regulator to balance exploration and exploitation. However, for online RL agents that require continuous environment interaction, the computational efficiency of these methods during inference becomes a critical concern. Score-matching-based methods typically require 20 diffusion steps, sometimes significantly more, during inference. Yet, when acceleration techniques like DPM-solver lu2022dpm are employed, reducing the number of steps below five still frequently leads to performance degradation. Policy gradient methods enable end-to-end optimization of the diffusion process, potentially approximating the target Boltzmann distribution using fewer denoising steps. However, the lack of strong intermediate supervision can lead to local optima and mode collapse.
In this work, we present DACER2: a significant step toward achieving stronger performance with fewer diffusion steps. The key contributions of this paper are the following: 1) We propose a Q-gradient field objective as an auxiliary optimization target to enhance diffusion policy training at each diffusion time step. 2) We propose a time-weight function $w(t)$ that takes the current diffusion time step as input to alleviate the problem that the Q-gradient field is independent of the diffusion time. This approach aligns with the denoising process requirements: higher amplitudes during early stages and precise control for smaller amplitudes in later stages. 3) We evaluated the efficiency and generality of our method on the popular MuJoCo benchmark. Compared with DACER wang2024diffusion , QVPO ding2024diffusion , QSM psenka2023learning , DIPO yang2023diffusion , DSAC duan2023dsac , TD3 fujimoto2018addressing , PPO schulman2017proximal , and SAC haarnoja2018soft , our approach achieved state-of-the-art (SOTA) performance. 4) We evaluated the training and inference times of all diffusion-based algorithms under identical hardware configurations using the PyTorch framework. Our method achieved 60.6% faster inference and 41.7% faster training compared to DACER, while also demonstrating the fastest inference efficiency. Detailed results are presented in Fig. 4.
2 Preliminaries
2.1 Reinforcement Learning with Soft Policy
RL problems are commonly modeled as Markov Decision Processes (MDPs) sutton2018reinforcement . An infinite-horizon MDP is defined by a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ is the state space and $\mathcal{A}$ is the action space, both assumed bounded and potentially continuous. $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ denotes the transition dynamics, specifying the probability distribution $P(\cdot\mid s_{t},a_{t})$ over next states, with $\Delta(\mathcal{S})$ representing the set of distributions over $\mathcal{S}$. $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, and $\gamma\in[0,1)$ is the discount factor.
The objective in RL is to find a policy $\pi$ that maximizes the expected discounted cumulative reward:
$J_{\pi}=\mathbb{E}_{(s_{i\geq t},a_{i\geq t})\sim\pi}\Big{[}\sum_{i=t}^{\infty}\gamma^{i-t}r(s_{i},a_{i})\Big{]},$ (1)
The state-action value function, or Q-function, for a policy $\pi$ is defined as:
$Q(s,a)=\mathbb{E}_{\pi}\Big{[}\sum_{i=0}^{\infty}\gamma^{i}r(s_{i},a_{i})\mid s_{0}=s,a_{0}=a\Big{]}.$ (2)
A central challenge in online RL is balancing exploration and exploitation to optimize long-term performance. One compelling strategy involves learning a policy that aims to approximate a soft policy haarnoja2017reinforcement ; haarnoja2018soft ; duan2021distributional ; duan2023dsac ; psenka2023learning ; ma2025soft ; messaoud2024s . Such target soft policies are typically formulated as a Boltzmann distribution, where the desired policy distribution is proportional to the exponentiated state-action value function:
$\pi_{\text{soft}}(a|s)\propto\exp\left(\frac{1}{\alpha}Q(s,a)\right).$ (3)
In the context of maximum entropy RL, this $Q(s,a)$ is often extended to a soft Q-function, denoted $Q_{\text{soft}}(s,a)$, which incorporates an entropy bonus into the standard value function to explicitly encourage exploration. The temperature parameter $\alpha>0$ controls the stochasticity of this target soft policy. The learned policy, $\pi$, is then trained to match this implicitly defined $\pi_{\text{soft}}$.
2.2 Diffusion Models as Expressive Policy
Diffusion models ho2020denoising ; song2020score ; wang2024diffusion conceptualize data generation as a stochastic process where data samples are iteratively reconstructed via a parameterized reverse-time stochastic differential equation (SDE). Although both forward and reverse diffusion processes are theoretically integral to these models, recent work by He et al. chen2024deconstructing highlights that their expressive power primarily stems from the reverse-time denoising dynamics, rather than the forward-time noising process. Accordingly, our analysis and modeling efforts concentrate on the reverse diffusion process.
Formally, the continuous reverse-time SDE that governs this process is defined as follows:
$d\boldsymbol{x}=\left[f(\boldsymbol{x},t)-g(t)^{2}\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})\right]dt+g(t)\,d\omega(t),$ (4)
where $f(\boldsymbol{x},t)$ represents the drift term, $g(t)$ denotes the time-dependent diffusion coefficient, $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$ is the score function, and $d\omega(t)$ is the standard Wiener process. The term $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$, also known as the score function, plays a central role in guiding the reverse diffusion dynamics. It is important to note that this equation represents the general form of the reverse-time SDE; the specific construction of terms such as $f(\boldsymbol{x},t)$ and $g(t)$ can vary across different diffusion model algorithms.
Therefore, the reverse-time SDE of diffusion policy can be expressed as:
$d\boldsymbol{a}_{t}=\left[f(\boldsymbol{a}_{t},t)-g(t)^{2}S_{\theta}(\boldsymbol{s},\boldsymbol{a}_{t},t)\right]dt+g(t)\,d\omega(t),$ (5)
where $S_{\theta}(\boldsymbol{s},\boldsymbol{a}_{t},t)$ is a neural network designed to approximate the gradient $\nabla_{\boldsymbol{a}_{t}}\log p_{t}(\boldsymbol{a}_{t}|\boldsymbol{s})$. Actions can be sampled from the diffusion policy $\pi_{\theta}(\boldsymbol{a}_{0}|\boldsymbol{s})$ by solving the following integral:
[…]
where $\boldsymbol{a}_{T}$ follows a standard normal distribution $\mathcal{N}(0,\boldsymbol{I})$.
2.3 Langevin Dynamics
Langevin dynamics represents a powerful computational framework for simulating particle motion under the joint influence of deterministic forces and stochastic fluctuations. When coupled with stochastic gradient descent, this approach gives rise to stochastic gradient Langevin dynamics (SGLD) welling2011bayesian - an efficient sampling algorithm that leverages log-probability gradients $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x})$ to draw samples from probability distributions $p(x)$ through an iterative Markov chain process:
$\boldsymbol{x}_{t-1}=\boldsymbol{x}_{t}+\frac{\delta_{t}}{2}\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{x}_{t})+\sqrt{\delta}_{t}\boldsymbol{\epsilon},$ (7)
where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $\delta_{t}$ is the step size. When $t$ range from infinity to one, $\delta_{t}\to 0$, $x_{0}$ equals to the true probability density $p(x)$.
3 Method
3.1 Q-Gradient Field Guided Denoising
Due to the intractability of computing the entropy in reverse SDEs, SDE-based policies face fundamental challenges when integrated with maximum entropy RL frameworks. To address this limitation, we shift our focus to soft policy. A key insight for SDE-based policy design is that policy improvement can be effectively achieved by directly maximizing the expected Q-value. The following theorem formalizes this principle, showing that policies which maximize the expected Q-value under a global entropy constraint inherently exhibit the structure of a soft policy.
Theorem 1.
Let $\mathcal{S}$ denote the state space and $\mathcal{A}$ denote the continuous action space. Suppose $p(s)$ is a distribution over states, $\mathcal{H}_{0}^{global}$ denotes a specific entropy value. We define the policy space $\Pi_{\mathcal{H}_{0}^{global}}$ as the set of policy families $\{\pi^{*}(\cdot|s)\}_{s\in\mathcal{S}}$, where each $\pi(\cdot|s)$ represents a valid probability distribution over actions. This policy family is required to satisfy a global expected entropy constraint:
$\mathbb{E}_{s\sim p(s)}[H(\pi^{*}(\cdot|s))]=\mathcal{H}_{0}^{global},$ (8)
where $\mathcal{H}_{0}^{global}$ is a given constant.
Within the policy space $\Pi_{\mathcal{H}_{0}^{global}}$, the family of policies $\{\pi^{*}(\cdot|s)\}_{s\in\mathcal{S}}$ that maximizes the global expected action value $\mathbb{E}_{s\sim p(s)}[\mathbb{E}_{a\sim\pi(a|s)}[Q(s,a)]$ has the property that, for each state $s$, the optimal policy $\pi^{*}(a|s)$ takes the form of a soft policy:
[…]
where $\alpha>0$ is a global temperature parameter, whose value is implicitly determined by a global expected entropy constraint: $\mathbb{E}_{s\sim p(s)}[H(\pi^{*}(\cdot|s))]=\mathcal{H}_{0}^{global}$.
Proof. See Appendix A.
Consequently, Theorem 1 inspires a novel design for SDE-based policy functions. By focusing on maximizing the expected Q-value, these policies can effectively learn a soft policy distribution.
Meanwhile, langevin dynamics provides an efficient method for sampling actions from Boltzmann distributions hinton2002training :
$\pi(\boldsymbol{a}|\boldsymbol{s})=\frac{e^{\frac{1}{\alpha}Q(\boldsymbol{s},\boldsymbol{a})}}{Z(\boldsymbol{s})},$ (10)
where $\alpha>0$ is temperature coefficient, $Q(\boldsymbol{s},\boldsymbol{a})$ is state-action value function and $Z(\boldsymbol{s})$ is the partition function that normalizes the distribution. The formula derived by taking the partial derivative of both sides of Eq. (10) with respect to $\boldsymbol{a}$ can be expressed as
$\nabla_{\boldsymbol{a}}\log\pi(\boldsymbol{a}|\boldsymbol{s})=\frac{1}{\alpha}\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a}).$ (11)
Substituting Eq. (11) into Eq. (7), we can obtain the sampling process for $\pi(\boldsymbol{a}|\boldsymbol{s})$:
$\boldsymbol{a}_{t-1}=\boldsymbol{a}_{t}+\frac{\delta_{t}}{2\alpha}\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})+\sqrt{\delta}_{t}\boldsymbol{\epsilon}.$ (12)
Equation (12) can be regarded as a trivial solution to the SDE-based policy that satisfies Theorem 1. This connection suggests using $\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})$ as an auxiliary learning objective to guide the training of the SDE-based policy, helping to ensure the policy avoids sharp local optima and maintains representational diversity to prevent mode collapse. The reason we do not rely solely on $\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})$ to train the score function is that training based exclusively on this gradient can lead to instability. Even minor noise in the $Q$-values may prevent the algorithm from converging to the optimal policy ding2024diffusion . Therefore, we use it only as an auxiliary training objective.
3.2 Time-weighted $\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})$
In the reverse-time SDE framework, directly maximizing the expected Q-value yields an entire family of valid solutions. However, the lack of strong intermediate supervision can lead to suboptimal solutions, such as local optima or mode collapse. To address this, we incorporate the gradient term $\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})$—which corresponds to Langevin dynamics, the most trivial member of this solution family—as an auxiliary objective when training the score network. However, the main issue is that $\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})$ is independent of the diffusion time step, whereas the score function is not. This time invariance cannot meet the denoising requirements that change throughout the diffusion process. Specifically, in the later stages of diffusion, the denoising intensity should naturally decrease as the action distribution approaches the optimal policy.
To address this issue, we introduce a time-dependent weight that modulates the influence of Q-gradient guidance based on the diffusion time step, allowing for more precise control over the denoising process. Inspired by the design approach for the step size $\delta_{t}$ in Eq. (7), we can similarly design $w(t)$ using the commonly employed exponential decay function welling2011bayesian ; teh2016consistency :
$w(t)=\exp(c\cdot t+d),$ (13)
where $t$ denotes the current diffusion step; following common practice, we keep the hyperparameters $c$ and $d$ fixed at 0.4 and -1.8 across all experiments.
Furthermore, during the training process, the instability in the estimation of the Q function leads to fluctuations in the numerical value of $\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})$, which in turn affects the stability of the training. To mitigate this issue, we normalize it by dividing its norm:
$\nabla_{\boldsymbol{a}_{t}}Q_{\text{norm}}(\boldsymbol{s},\boldsymbol{a}_{t})=\frac{\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})}{||\nabla_{\boldsymbol{a}_{t}}Q(\boldsymbol{s},\boldsymbol{a}_{t})||+\epsilon},$ (14)
where $\epsilon$ is a small constant to prevent division by zero.
Ultimately, we construct the Q-gradient field objective function to facilitate the training of the diffusion policy, where $\pi_{\theta}(\boldsymbol{a}_{t}|\boldsymbol{s})$ denotes the action generated using the diffusion policy as defined in Eq. (6):
$\mathcal{L}_{g}(\theta)=\min_{\theta}\underset{\begin{subarray}{c}\boldsymbol{s}\sim\mathcal{B}\\ t\sim\text{U}(1,T)\\ \boldsymbol{a}_{t}\sim\pi_{\theta}(\boldsymbol{a}_{t}|\boldsymbol{s})\end{subarray}}{\mathbb{E}}\left[\left\|w(t)\nabla_{\boldsymbol{a}_{t}}Q_{\text{norm}}(\boldsymbol{s},\boldsymbol{a}_{t})-S_{\theta}(\boldsymbol{s},\boldsymbol{a}_{t},t)\right\|_{2}^{2}\right],$ (15)
where U means uniform distribution, $t$ is the current diffusion step, $\mathcal{B}$ represents the replay buffer, and $\theta$ is the network parameter of the diffusion policy. The subscript $g$ represents the objective function related to the Q-gradient.
3.3 DACER2: A High Efficiency Diffusion RL Algorithm
In the critic component, we adopt the double Q-learning strategy fujimoto2018addressing to alleviate overestimation bias. Specifically, we maintain two independent Q-function estimators, denoted as $Q_{\phi_{1}}(\boldsymbol{s},\boldsymbol{a})$ and $Q_{\phi_{2}}(\boldsymbol{s},\boldsymbol{a})$, which are trained to approximate the true action-value function. To enhance training stability, we introduce two corresponding target networks, $Q_{\bar{\phi}_{1}}(\boldsymbol{s},\boldsymbol{a})$ and $Q_{\bar{\phi}_{2}}(\boldsymbol{s},\boldsymbol{a})$, which are updated softly from the main networks following the technique in van2016deep .
The Q-networks are optimized by minimizing the Bellman error. For each network $Q_{\phi_{i}}(\boldsymbol{s},\boldsymbol{a})$, the loss $J_{Q}(\phi_{i})$ is defined as:
$J_{Q}(\phi_{i})\!=\!\underset{\begin{subarray}{c}(\boldsymbol{s},\boldsymbol{a},\boldsymbol{r},\boldsymbol{s}^{\prime})\sim\mathcal{B}\\ \boldsymbol{a}^{\prime}\sim\pi_{\theta}(\boldsymbol{a}_{0}|\boldsymbol{s})\end{subarray}}{\mathbb{E}}\left[\!\left(\!\left(r(\boldsymbol{s},\boldsymbol{a})\!+\!\gamma\min_{i=1,2}Q_{\bar{\phi}_{i}}(\boldsymbol{s}^{\prime},\boldsymbol{a}^{\prime})\!\right)-\!Q_{\phi_{i}}(\boldsymbol{s},\boldsymbol{a})\!\right)^{2}\right],\$ (16)
where $\gamma$ is discount factor, the target is computed as the smaller of the two target Q-values, $Q_{\bar{\phi}_{1}}(\boldsymbol{s}^{\prime},\boldsymbol{a}^{\prime})$ and $Q_{\bar{\phi}_{2}}(\boldsymbol{s}^{\prime},\boldsymbol{a}^{\prime})$, to prevent over-optimistic estimates. Furthermore, we incorporate the distributional value estimation framework from DSAC duan2023dsac to further mitigate overestimation issues.
In the actor component, we follow the objective function of maximizing the Q value and combine it with the auxiliary learning objective based on the Q-gradient field proposed in this paper. The final policy-learning objective is a linear combination:
$\pi=\arg\min_{\pi_{\theta}}\mathcal{L_{\pi}}(\theta)=\mathcal{L}_{q}(\theta)+\eta\cdot\mathcal{L}_{g}(\theta),$ (17)
[…]
where $\eta$ is a hyperparameter, $\mathcal{L}_{q}(\theta)=\mathbb{E}_{\boldsymbol{s}\sim\mathcal{B},\boldsymbol{a}_{0}\sim\pi_{\theta}(\cdot|\boldsymbol{s})}\left[-Q_{\phi}(\boldsymbol{s},\boldsymbol{a}_{0})\right],$ $p(s)$ is a distribution over states.
4 Experimental Results
This section presents the experimental results. We first evaluate the multimodality of DACER2 and DACER using a toy environment called "Multi-goal" haarnoja2017reinforcement , as illustrated in Fig. 2. Then we show the empirical results of the proposed DACER2 algorithm evaluated with OpenAI Gym MuJoCo 2012MuJoCo tasks. With these experimental results, we aim to answer three questions:
  • Does DACER2 demonstrate stronger multimodal capabilities compared to DACER?
  • How does DACER2 compare to previous popular online RL algorithms and existing diffusion-based online RL algorithms?
  • How does the inference and training efficiency of DACER2 compare with existing diffusion-based RL methods?
Baselines.
The baselines encompass two categories of model-free reinforcement learning algorithms. The first category consists of diffusion-based RL methods, including a range of recent diffusion-policy online algorithms such as DACER wang2024diffusion , QVPO ding2024diffusion , DIPO yang2023diffusion , and QSM psenka2023learning . The second category includes classic model-free online RL baselines, namely TD3 fujimoto2018addressing , SAC haarnoja2018soft , PPO schulman2017proximal , and DSAC duan2023dsac .
Evaluation Setups.
We implemented our algorithm in JAX and evaluated it on eight MuJoCo tasks using the same metrics as DACER. Experiments were conducted on a system equipped with an AMD Ryzen Threadripper 3960X 24-core processor and an NVIDIA GeForce RTX 4090 GPU. For example, training Walker2d for 1.5M steps takes about two hours. Performance was measured by averaging the top returns from the final 10% of iterations over ten episodes. Results across five random seeds are reported with mean and standard deviation. For classic model-free baselines, we cite DACER-reported results, while all diffusion-based methods were re-evaluated. Furthermore, the training curves presented in Fig. 3 demonstrate the stability of the training process.
4.1 Toy Example
We evaluate the trained policy in the "Multi-goal" environment by initializing the agent at the origin and sampling 100 trajectories. The agent is tasked with navigating toward one of six symmetrically arranged goal points. As illustrated in Fig. 2, DACER2 demonstrates significantly enhanced multimodality compared to DACER. Specifically, DACER2 consistently reaches all six target locations with approximately uniform coverage, indicating that it effectively captures the underlying multimodal structure of the task and maintains balanced expression across distinct behavioral modes.
In contrast, DACER only reaches five out of the six goals, and its visitation frequencies are noticeably uneven. This imbalance suggests mode collapse and reveals limitations in its ability to represent diverse behaviors. These results highlight that DACER2 is better equipped to express diverse, mode-separated policies in multimodal environments. Additional multimodal experiments are provided in Appendix C.
(a) DACER2
(b) DACER
Figure 2: Multi-goal Task. Trajectories generated by a policy learned using our method (solid blue lines) are shown, with the $x$-axis and $y$-axis representing 2D positions (states). The agent is initialized at the origin, and the goals are marked as red dots. The level curves indicate the reward, and reaching within 1 of the endpoint signifies task completion. On the left are the experimental results of DACER2, and on the right are the experimental results of DACER.
4.2 Comparative Evaluation
The performance comparison, as summarized in Table 1 and illustrated in Fig. 3, indicates that our method consistently outperforms all baseline algorithms across the OpenAI Gym MuJoCo benchmark tasks. In particular, in more challenging locomotion environments such as Humanoid, Ant, HalfCheetah, and Walker2d, our method achieves substantial improvements over SAC, with respective gains of 27.4%, 43.4%, 9.3%, and 24.1%.
Furthermore, among diffusion-based online RL approaches, DACER2 further improves upon DACER, achieving performance gains of 0.2%, 15.3%, 3.7%, and 16.4% in the same set of tasks. Notably, DACER2 attains performance that is comparable to or even surpasses DACER, while requiring only five diffusion steps. These results collectively demonstrate the robustness and efficiency of our method, and further underscore the potential of diffusion-based policies for online RL.
(a) Humanoid-v3
(b) Ant-v3
(c) HalfCheetah-v3
(d) Walker2d-v3
(e) Inverted2Pendulum-v3
(f) Hopper-v3
(g) Pusher-v2
(h) Swimmer-v3
Figure 3: Training curves on benchmarks. The solid lines represent the mean, while the shaded regions indicate the 95% confidence interval over five runs. For PPO, iterations are defined by the number of network updates.
Table 1: Average final return. Computed as the mean of the highest return values observed in the final 10% of iteration steps per run. The value for each task is bolded. ± corresponds to standard deviation over five runs.
Humanoid Ant HalfCheetah Walker2d
Classic Model-Free RL PPO $6869\pm 1563$ $6157\pm 185$ $5789\pm 2201$ $4832\pm 638$
TD3 $5632\pm 436$ $6184\pm 487$ $8633\pm 4041$ $5237\pm 336$
SAC $9335\pm 696$ $6427\pm 804$ $16573\pm 224$ $6201\pm 263$
DSAC $10829\pm 243$ $7086\pm 261$ $17025\pm 157$ $6424\pm 147$
Diffusion Policy RL QSM $6072\pm 691$ $4783\pm 1235$ $11401\pm 882$ $5685\pm 437$
DIPO $9353\pm 356$ $3449\pm 149$ $12267\pm 2180$ $5066\pm 365$
DACER $11868\pm 56$ $7994\pm 61$ $17466\pm 96$ $6610\pm 12$
QVPO $9656\pm 252$ $6484\pm 145$ $14355\pm 175$ $6057\pm 352$
DACER2 (ours) $\mathbf{11896\pm 53}$ $\mathbf{9217\pm 64}$ $\mathbf{18107\pm 146}$ $\mathbf{7694\pm 352}$
IDP Hopper Pusher Swimmer
Classic Model-Free RL PPO $9357\pm 2$ $2647\pm 481$ $-22.9\pm 1.4$ $130\pm 2$
TD3 $9347\pm 15$ $3569\pm 455$ $-21.4\pm 1.2$ $134\pm 5$
SAC $\mathbf{9360\pm 0}$ $2483\pm 943$ $-19.6\pm 0.3$ $140\pm 14$
DSAC $\mathbf{9360\pm 0}$ $3660\pm 533$ $\mathbf{-19.4\pm 0.9}$ $138\pm 6$
Diffusion Policy RL QSM $591\pm 98$ $2006\pm 251$ $-73.8\pm 3.0$ $46\pm 1$
DIPO $9355\pm 2$ $3839\pm 40$ $-20.7\pm 0.2$ $55\pm 2$
DACER $9354\pm 2$ $4094\pm 54$ $-19.7\pm 0.2$ $150\pm 4$
QVPO $9354\pm 5$ $4046\pm 94$ $-20.4\pm 0.1$ $130\pm 10$
DACER2 (ours) $9359\pm 1$ $\mathbf{4213\pm 91}$ $\mathbf{-19.4\pm 0.1}$ $\mathbf{165\pm 2}$
4.3 Efficiency Analysis
We evaluate the training and inference efficiency of the algorithm, as illustrated in Fig. 4. All algorithms are implemented and tested using PyTorch to reflect practical deployment conditions, as PyTorch is commonly adopted in real-world control systems for both training and inference. Compared to the three diffusion-based algorithms, DACER2 is 41.7%, 49.3%, and 50.0% faster than DACER, QVPO, and DIPO, respectively, but 16.7% slower than QSM in terms of training speed. Given that the performance of our algorithm significantly outperforms QSM, the minor difference in training speed can be disregarded. Regarding inference speed, DACER2 shows a 60.6%, 82.5%, 60.6%, and 60.6% improvement over DACER, QVPO, QSM, and DIPO, respectively.
DACER2 adopts a Q-gradient field to assist the score function training under the max-Q objective framework, allowing more effective guidance during each denoising step. This design improves sample efficiency and accelerates convergence without requiring a large number of diffusion steps. As a result, even with only five diffusion steps, our algorithm achieves strong performance and substantial gains in training and inference efficiency. In real-time industrial control tasks, the inference time should be less than 1 millisecond to meet control requirements. Among the existing diffusion-based algorithms, only DACER2 meets this constraint.
Figure 4: Comparison of inference and training time. The left subfigure depicts the time required for each algorithm to complete a training step, excluding the time spent interacting with the environment. The right subfigure depicts the time required for the policy network to output an action after receiving a single state as input. All results are based on five different seeds.
4.4 Ablation Study
In this section, we conduct ablation study to investigate the impact of the following two aspects on the performance of the diffusion policy: 1) whether to use the Q-gradient field training objective function derived from the connection between noise prediction network and Q-gradient field; 2) whether to weight $\nabla_{a}Q(s,a)$ according to the time step of the diffusion process.
(a) Ablation for the Q-gradient field training objective function.
(b) Ablation for the time-weighted $\nabla_{a}Q(s,a)$.
Figure 5: Ablation experiment curves. (a) DACER2 represents the algorithm with Q-gradient field loss, and 20 steps represents the diffusion step size of 20. DACER2’s performance on Walker2d-v3 is far better than DACER. (b) Time-weighted means using $w(t)\nabla_{a}Q(s,a)$ instead of $\nabla_{a}Q(s,a)$ as the target value in the Q-gradient field training loss.
Q-gradient field training objective function.
In this ablation study, we first fixed the diffusion step size at 5 to evaluate the impact of the Q-gradient field loss function. As shown in Fig. 5(a), removing this objective—yielding a variant equivalent to DACER with 5 steps—led to a significant decline in performance. Next, we increased the diffusion step size of the algorithm without the Q-gradient field loss to 20; however, its performance still remained inferior to that of DACER2 with only 5 diffusion steps. These results demonstrate that the Q-gradient field loss function plays a crucial role in guiding the diffusion denoising process and serves as a key component for enhancing overall performance.
Time-weighted $\nabla_{a}Q(s,a)$.
We conducted an experiment to demonstrate that using time-weighted $\nabla_{a}Q(s,a)$ can further improve performance. As shown in Fig. 5(b), directly using $\nabla_{a}Q(s,a)$ as the target value in the Q-gradient field training loss, instead of the $w(t)\nabla_{a}Q(s,a)$, results in performance degradation. This is because different timesteps require matching different magnitudes of noise prediction, which enhances both training stability and final performance.
5 Related Work
In this section, we review existing works on using the diffusion model as a policy function in combination with RL.
Online RL with Diffusion Policy.
Online RL enables agents to refine their policies through real-time interaction. Yang et al. introduced DIPO yang2023policy , which maintains a dedicated diffusion buffer to store actions and model them using diffusion techniques. Psenka et al. proposed QSM psenka2023learning , which aligns policies with $\nabla_{a}Q$ via score matching, but is sensitive to value gradient inaccuracies across the action space. Recently, Ding et al. ding2024diffusion proposed QVPO, which weights diffusion-sampled actions by Q-values without computing gradients. However, it uses a fixed ratio of uniform samples to boost the entropy, lacking adaptive control and later degrading performance. Ma et al. ma2025soft proposed SDAC, which uses score matching over noisy energy-based diffusion. It avoids requiring optimal actions but suffers from high gradient variance due to poor sampling in high-Q regions.
Offline RL with Diffusion Policy.
Offline RL focuses on learning optimal policies from suboptimal datasets, with the core challenge being the out-of-distribution (OOD) problem kumar2020conservative ; fujimoto2019off . Diffusion models are naturally suited for offline RL due to their ability to model complex data distributions. Wang et al. proposed Diffusion-QL wang2022diffusion , which combines behavior cloning through a diffusion loss with Q-learning to improve policy learning. However, Diffusion-QL suffers from slow training and instability in OOD regions. To address the former, Kang et al. proposed Efficient Diffusion Policy (EDP) kang2023efficient , which speeds up training by initializing from dataset actions and adopting a one-step sampling strategy. To mitigate OOD issues, Ada et al. introduced SRDP ada2024diffusion , which enhances generalization by integrating state reconstruction into the diffusion policy. Furthermore, Chen et al. proposed CPQL chen2023boosting , a consistency-based method that improves efficiency via one-step noise-to-action generation during both training and inference, albeit with some performance trade-offs.
Comparison with DACER.
Wang et al. proposed DACER wang2024diffusion , which uses the reverse diffusion process as a policy approximator and employs a Gaussian mixture model (GMM) to estimate policy entropy for balancing exploration and exploitation. However, the work lacks theoretical justification for why maximizing the expected Q-value enables learning multimodal policies. Moreover, a key trade-off remains: long diffusion processes hinder training efficiency, while fewer steps lead to performance degradation. Our work addresses this by reducing the diffusion steps while maintaining or even improving both performance and policy multimodality.
6 Conclusion
In this paper, we address the critical challenge of balancing inference efficiency and performance in diffusion-based online RL. By introducing a Q-gradient field objective and a time-dependent weighting scheme, our method enables each denoising step to be guided by the Q-function with adaptive emphasis over time. This design allows the policy to achieve strong performance with only five diffusion steps, significantly improving both training and inference speed. In addition, our framework inspires a novel class of SDE-based policy functions that maximize the expected Q-value under a fixed entropy constraint, enabling effective learning of expressive soft policy distributions.
Appendix A Theoretical Results
Proof. Let $\mathcal{S}$ be the state space, $\mathcal{A}$ denotes the continuous action space, and $p(s)$ is a probability measure over states. We seek a family of policies $\{\pi(\cdot\mid s)\}_{s\in\mathcal{S}}$ belonging to the constrained space:
$\Pi_{\mathcal{H}^{\text{global}}_{0}}\;=\;\Bigl{\{}\{\pi(\cdot\mid s)\}_{s\in\mathcal{S}}\ \Bigm{|}\ \mathbb{E}_{s\sim p(s)}\bigl{[}H(\pi(\cdot\mid s))\bigr{]}=\mathcal{H}^{\text{global}}_{0},\;\int_{\mathcal{A}}\!\pi(a\mid s)\,da=1,\ \forall s\Bigr{\}},$ (18)
which maximises the expected action-value
$J\bigl{(}\{\pi(\cdot\mid s)\}\bigr{)}\;=\;\mathbb{E}_{s\sim p(s)}\!\Bigl{[}\mathbb{E}_{a\sim\pi(\cdot\mid s)}\!\bigl{[}Q(s,a)\bigr{]}\Bigr{]}\;=\;\int_{\mathcal{S}}p(s)\!\!\int_{\mathcal{A}}\pi(a\mid s)\,Q(s,a)\,da\,ds.$ (19)
Then, we introduce a scalar multiplier $\alpha$ for the global expected-entropy constraint and a state-dependent multiplier $\eta(s)$ for the normalisation constraint at each $s$. The Lagrangian reads
[…]
Because the decision variables for distinct states couple only through $\alpha$, we can minimise the integrand for each fixed $s$ independently:
$\mathcal{L}_{s}\bigl{(}\pi(\cdot\mid s)\bigr{)}\;=\;\int_{\mathcal{A}}\Bigl{[}p(s)\pi(a\mid s)Q(s,a)-\alpha\,p(s)\pi(a\mid s)\log\pi(a\mid s)+\eta(s)\pi(a\mid s)\Bigr{]}\,da.$ (21)
Taking the functional derivative and setting it to zero yields, for almost every $a\in\mathcal{A}$, we can obtain
[…]
Assuming $p(s)>0$, we divide both sides by $p(s)$ and rearrange:
[…]
Let $\tilde{\eta}(s)=\eta(s)/p(s)$. Exponentiating gives the unnormalised form
$\pi(a\mid s)\;=\;\exp\!\Bigl{(}\frac{\tilde{\eta}(s)-\alpha}{\alpha}\Bigr{)}\exp\!\Bigl{(}\tfrac{Q(s,a)}{\alpha}\Bigr{)}\;=\;C(s)\,\exp\!\Bigl{(}\tfrac{Q(s,a)}{\alpha}\Bigr{)},$ (24)
where $C(s)$ is a state-wise normalising constant.
[…]
Therefore, the optimal policy family is the Boltzmann distribution
[…]
The scalar $\alpha>0$ is the Lagrange multiplier associated with the global entropy constraint and serves as a common temperature across all states. Its value is obtained implicitly by substituting $\pi^{*}$ back into
$\mathbb{E}_{s\sim p(s)}\!\bigl{[}H(\pi^{*}(\cdot\mid s))\bigr{]}=\mathcal{H}^{\text{global}}_{0}.$ (27)
Consequently, although the entropy constraint is imposed only on the state-averaged entropy, each per-state optimal policy still follows a Boltzmann form with the same temperature parameter $\alpha$.
Appendix B Environmental Details
MuJoCo [31]:
This is a high-performance physics simulation platform widely adopted for robotic reinforcement learning research. The environment features efficient physics computation, accurate dynamic system modeling, and comprehensive support for articulated robots, making it an ideal benchmark for RL algorithm development. In this research, we concentrate on eight tasks: Humanoid, Ant, HalfCheetah, Walker2d, InvertedDoublePendulum (IDP), Hopper, Pusher, and Swimmer, as illustrated in Fig. 6. The IDP task entails maintaining the balance of a double pendulum in an inverted state. In contrast, the objective of the other tasks is to maximize the forward velocity while avoiding falling. All these tasks are realized through the OpenAI Gym interface [2].
Figure 6: MuJoCo simulation tasks. (a) Humanoid-v3:$(s\times a)\in\mathbb{R}^{376}\times\mathbb{R}^{17}$. (b) Ant-v3: $(s\times a)\in\mathbb{R}^{111}\times\mathbb{R}^{8}$. (c) HalfCheetah-v3 : $(s\times a)\in\mathbb{R}^{17}\times\mathbb{R}^{6}$. (d) Walker2d-v3: $(s\times a)\in\mathbb{R}^{17}\times\mathbb{R}^{6}$. (e) InvertedDoublePendulum-v3: $(s\times a)\in\mathbb{R}^{6}\times\mathbb{R}^{1}$. (f) Hopper-v3: $(s\times a)\in\mathbb{R}^{11}\times\mathbb{R}^{3}$. (g) Pusher-v2: $(s\times a)\in\mathbb{R}^{23}\times\mathbb{R}^{7}$. (h) Swimmer-v3: $(s\times a)\in\mathbb{R}^{8}\times\mathbb{R}^{2}$.
Appendix C Additional multimodal experiments
The MuJoCo benchmarks also demonstrate multimodal action distributions across robotic control tasks, as shown in Fig. 7. By leveraging the trained DACER2 to conduct multiple action samplings in identical states, it can be observed that the policy manifests multiple behavioral modes in specific states. Each subfigure depicts 100 actions sampled for the respective state, projected into $\mathbb{R}^{2}$ using UMAP [22] for visualization.
We conducted the same experiment with DACER and observed that while multimodal behaviors are still present in some tasks, the Ant and Walker2d task fails to show clear multimodal patterns under the DACER policy. These results suggest that DACER2 exhibits stronger and more consistent multimodality than DACER across MuJoCo tasks.
Table 2: Hyperparameters used in DACER2.
Parameter Humanoid Ant HalfCheetah Walker2d
Loss function weight $\eta$ 2.0 2.0 2.0 1.0
Parameter IDP Hopper Pusher Swimmer
Loss function weight $\eta$ 0.1 0.1 0.1 0.001
Appendix D Experimental Hyperparameters
The hyperparameters of all baseline algorithms except the diffusion-based algorithm are shown in Table 3. Additionally, the parameters for all diffusion-based algorithms, including DACER2, are presented in Table 2 and Table 4.
Table 3: Baseline hyperparameters.
Hyperparameters Value
Shared
   Replay buffer capacity 2,000,000
   Buffer warm-up size 30,000
   Batch size 256
   Action bound $[-1,1]$
   Hidden layers in critic network [256, 256, 256]
   Hidden layers in actor network [256, 256, 256]
   Activation in critic network GeLU
   Activation in actor network GeLU
   Optimizer Adam ($\beta_{1}=0.9,\beta_{2}=0.999$)
   Actor learning rate $41{\rm{e-}}$ 4
   Critic learning rate $41{\rm{e-}}$ 4
   Discount factor ($\gamma$) 0.99
   Policy update interval 2
   Target smoothing coefficient ($\rho$) 0.005
   Reward scale 0.2
Maximum-entropy framework
   Learning rate of $\alpha$ $43{\rm{e-}}$ 4
   Expected entropy ($\overline{\mathcal{H}}$) ¯$\overline{\mathcal{H}}=-{\rm{dim}}(\mathcal{A})$
Deterministic policy
   Exploration noise $\epsilon\sim\mathcal{N}(0,0.1^{2})$
Off-policy
   Replay buffer size $1\times 10^{6}$
   Sample batch size 20
On-policy
   Sample batch size 2,000
   Replay batch size 2,000
Table 4: Diffusion-based algorithms’ hyperparameter
Parameter DACER2 DACER QVPO QSM DIPO
Replay buffer capacity 2e6 2e6 2e6 2e6 2e6
Buffer warm-up size 3e4 3e4 3e4 3e4 3e4
Batch size 256 256 256 256 256
Discount $\gamma$ 0.99 0.99 0.99 0.99 0.99
Target network soft-update rate $\rho$ 0.005 0.005 0.005 0.005 0.005
Network update times per iteration 1 1 1 1 1
Action bound $[-1,1]$ $[-1,1]$ $[-1,1]$ $[-1,1]$ $[-1,1]$
Reward scale 0.2 0.2 0.2 0.2 0.2
No. of hidden layers 2 2 2 2 2
No. of hidden nodes 256 256 256 256 256
Activations in critic network GeLU GeLU Mish ReLU Mish
Activations in actor network Mish Mish Mish ReLU Mish
Diffusion steps 5 20 20 20 20
Policy delay update 2 2 2 2 2
Action gradient steps N/A N/A N/A N/A 20
No. of Gaussian distributions 3 3 N/A N/A N/A
No. of action samples 200 200 N/A N/A N/A
Time-weighted hyperparameter $c$ 0.4 N/A N/A N/A N/A
Time-weighted hyperparameter $d$1.8 N/A N/A N/A N/A
Alpha delay update 10,000 10,000 N/A N/A N/A
Noise scale $\lambda$ 0.1 0.1 N/A N/A N/A
Optimizer Adam Adam Adam Adam Adam
Actor learning rate $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$
Critic learning rate $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$
Alpha learning rate $3\cdot 10^{-2}$ $3\cdot 10^{-2}$ N/A N/A N/A
Target entropy $-{\rm{dim}}(\mathcal{A})$ $-{\rm{dim}}(\mathcal{A})$ N/A N/A N/A
Figure 7: Demonstration of multimodal action distributions on MuJoCo benchmarks. From top to bottom they are Humanoid-v3, Walker2d-v3, HalfCheetah-v3, Ant-v3. The figure displays sampled actions from DACER2. Each subfigure depicts 100 actions sampled for the respective state, projected into $\mathbb{R}^{2}$ using UMAP [22].
Appendix E Limitation and Future Work
In this study, we propose the Q-gradient field objective as an auxiliary training loss to guide the diffusion policy with more informative gradient signals. Although we employ a normalization technique on $\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})$ to enhance training stability, the errors and instability inherent in Q-value estimation can still adversely affect the training process. In future work, we plan to incorporate techniques such as ensemble Q-learning and multimodal value function estimation to mitigate Q-value estimation errors and instability, thereby further improving the robustness of our algorithm.